Sally travels 460m south 980m west find the displacement and distance?
1
Expert's answer
2020-03-11T10:29:50-0400

Distance:


"d=460+980=1440\ m"

Displacement:


"s=\sqrt{460^2+980^2}=1083\ m"

Direction:


"\theta=\arctan{\frac{460}{980}}=25.1\degree\text{S of W}"